OpenStudy (anonymous):

x4 + 12x2 - 64 i know the answer is (x2 - 4)(x2 + 16) just don't know how

OpenStudy (anonymous):

Let y=x^2 x^4+12x^2-64=y^2+12y-64. This you can factorise to get (y-4)(y+16) Then just replace y with x^2 again.

OpenStudy (anonymous):

moorsy is correct, if you understand how to factorize quadratics, then the one above is simple.

OpenStudy (anonymous):

If you do not understand how to factorize quadratics, then that is another problem

OpenStudy (anonymous):

me, understanding, math -all those words in the same sentence don't work ;) but i am slowly getting it, thank you!